# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v1

@@ 0.D.3.3 Results
Experiment 1 10000 + $R_{E}$ km 0.1 $5.0^{\circ}$ $10.0^{\circ}$ $10.0^{\circ}$ $10.0^{\circ}$
Experiment 2 15000 + $R_{E}$ km 0.5 $8.0^{\circ}$ $10.0^{\circ}$ $10.0^{\circ}$ $90.0^{\circ}$
Experiment 3 5000 + $R_{E}$ km 0.001 $5.0^{\circ}$ $10.0^{\circ}$ $180.0^{\circ}$ $180.0^{\circ}$
@@
Figure 16: Reward (or error) per step, for each experiment. An error of 0.0 indicates that the follower contains exactly the same orbital elements as the leader, including anomaly.
@@
This experiment demonstrates that the agent learns an effective policy capable of generalizing thrust maneuvers to initial conditions it did not encounter during training, while also adapting to realistic orbital dynamics with non-conservative forces. Future improvements to this mission could include introducing additional penalties for fuel consumption in the reward function and further exploring the limits of the agent’s strategy.
@@
0.D.4 Collision Avoidance
0.D.4.1 Definition
@@
where $\hat{i}$, $\hat{j}$ and $\hat{k}$ are the unit vectors forming the orthogonal system. Given the assumption of negligible velocity uncertainty, only the $\hat{j}$ and $\hat{k}$ axes (defining the conjunction plane) are relevant for the analysis. This simplifies the problem to two dimensions by introducing a projection matrix $C=[\hat{j},\hat{k}]$.
Under the assumption that uncertainties follow a Gaussian distribution, the distribution’s parameters are given by $\mu=Cr_{\text{RTN}}$ and $\Sigma=C(\Sigma_{\text{T,RTN}}+\Sigma_{\text{C,RTN}})C^{T}$. The PoC is then calculated as the integral over a circle centered at the origin with radius $R=R_{T}+R_{C}$:
[…]
@@
In this approach, we use a faster computation method compared to the traditional one, as outlined in [20]. A PoC value greater than $10^{-6}$ indicates a high collision risk, requiring maneuvers to address it.
0.D.4.2 Environment Setup
@@
The observation $o_{t}=(s_{t},M_{t},s^{\prime}_{t},M^{\prime}_{t},f_{t},P_{t})\in\mathbb{R}^{14}$ consists of the current equinoctial elements of the satellite ($s_{t}$) and the drifter ($s^{\prime}_{t}$), both bodies mean anomaly ($M_{t}$ and $M^{\prime}_{t}$), together with the current fuel ($f_{t}$) and PoC ($P_{t}$).
As in the Hohmann maneuver mission (0.D.2), mission, apart from the polar representation of the thrust, ithe action space also contains the decision ($\delta$) to perform the given thrust: $a_{t}=(T,\theta,\phi,\delta)$. This action space is limited by $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}},\delta_{\text{max}})=(5,\pi,2\pi,1)$.
@@
where $\Delta s_{t}=w^{T}(s_{t}-\hat{s})$, with $w^{T}$ representing a vector of weights, $P_{t-1}$ and $P_{t}$ are respectively the PoC before and after the action, and $\alpha_{1}$ and $\alpha_{2}$ are weights given to each parameter.
0.D.4.3 Results
The training hyperparameters for the agent are listed in Table 9. To assess the agent’s performance, we analyze its decision-making process based on the simulated PoC, miss distance, and TCA. Additionally, we evaluate the agent’s generalization ability by enabling all perturbative forces available in OrbitZoo, including harmonic gravity fields, third body forces (Sun and Moon), solar radiation pressure, and drag. However, the satellite simulator dynamics remain unchanged, maintaining the realistic constraint of incomplete information available to the agent.
Table 9: Collision Avoidance: Training hyperparameters.
@@ 0.D.4.3 Results
After training, Figure 18 demonstrates that the agent applies thrust early in the episode ($\delta>0.5$) for both unrealistic and realistic dynamics, successfully reducing the simulated PoC for the remainder of the episode. However, when using realistic dynamics, the simulated PoC becomes more unstable, leading to greater uncertainty for the agent when deciding when and how to apply thrust.
When the agent is given 4 days to execute the maneuvers, as shown in Figure 19, a similar strategy emerges. However, the PoC at the initially detected TCA increases unexpectedly, indicating that the simulator’s values were insufficient to predict this behavior accurately. By analyzing the changes in orbital elements, as depicted in Figure 20, we see that the agent’s strategy involves increasing the semi-major axis by 3 km while keeping the eccentricity close to its nominal value. Although the agent has the capability to make adjustments in the cross-track direction, it correctly concludes that changing the inclination is unnecessary to minimize the collision risk. A comparable policy is observed when realistic dynamics are applied.
From these experiments, we conclude that the agent can generalize its policy to realistic dynamics when given the same amount of time (2 days) to perform the maneuvers as during training. However, when the agent is provided with 4 days to take action, the uncertainty in the positions of both bodies negatively impacts the policy, leading to an unexpected increase in the PoC at the initial TCA. Future improvements could involve training the agent with a random number of days to complete the maneuvers and incorporating more informed observations to reduce uncertainty and improve decision-making.
@@
Figure 18: Collision Avoidance in 2 days: Agent decision to apply thrust ($\delta>0.5$), Probability of Collision (POC), Miss Distance, and Time of Closest Approach (TCA) by step.
Figure 19: Collision Avoidance in 4 days: Agent decision to apply thrust ($\delta>0.5$), Probability of Collision (POC), Miss Distance, and Time of Closest Approach (TCA) by step.
@@
Figure 20: Evolution of the semi-major axis (a) and eccentricity components ($e_{x}$ and $e_{y}$) of the orbit by step, in the 2 days and 4 days approach.
0.D.5 GEO Constellation
0.D.5.1 Definition
@@
There are several possible types of formation that constellations of satellites in GEO can have. Here, we focus on a mission in which all agents learn to stay at equal distances around the orbit.
0.D.5.2 Environment Setup
We define an environment with $n=4$ agents in GEO, characterized by a semi-major axis of $a_{\text{GEO}}=35786+R_{E}=42164$ km, where $R_{E}$ is Earth’s equatorial radius. Each agent is a satellite with $m_{0}=250$ kg, 50 kg of which is fuel, and equipped with a thrust with $I_{sp}=3100$ s.
Each episode contains 2000 steps with step sizes of 360 s, and the dynamics are purely based on Newtonian attraction. At the start of each episode ($t=0$), agents are initialized with random mean anomalies ($M_{0}^{i},i=\{1,2,...,n\}$), representing their positions along the orbit. The objective for the agents is to learn to distribute themselves uniformly around the orbit, maintaining angular separations of $2\pi/n$ radians (90 degrees for $n=4$).
@@ 0.D.5.2 Environment Setup
The observation space is also reduced. Each agent observation receives its current semi-major axis, eccentricity components, remaining fuel, and the anomalies of all agents in the environment: $o_{t}=(a_{t},e_{xt},e_{yt},f_{t},M_{t}^{1},M_{t}^{2},M_{t}^{3},M_{t}^{4})$.
To calculate the collective penalty related to the difference in anomalies ($P_{M}$), we first normalize the anomalies to be within $[0,2\pi]$, followed by:
@@
0.D.5.3 Results
The agents were trained independently, without shared observations for the critic network training, with the training hyperparameters provided in Table 10. After training, we tested the generalization capability of the policy by conducting Monte Carlo simulations to analyze the trends in penalty reduction for the anomalies. These trends are shown in Figure 21, using both simple dynamics (as used during training) and realistic dynamics, which include harmonic gravity fields, third body forces, solar radiation pressure, and drag.
The results indicate that the agents effectively reduce the collective penalty (anomalies distance) in both scenarios. However, their specific ability to achieve this depends on the initial configuration of the constellation. Through the OrbitZoo interface, shown in Figure 5, it is evident that the agents are successfully learning the intended objective, albeit with some imperfections.
In conclusion, even though the agents were trained independently without any information sharing, they were able to effectively learn how to maintain equal distances from one another around the GEO orbit, even when using realistic dynamics that were not part of the training. Future improvements include implementing algorithms specifically designed for Multi-Agent Reinforcement Learning (MARL), such as MAPPO, enabling information sharing through Federated Reinforcement Learning (FRL), and scaling to larger constellations of bodies to represent more realistic scenarios.
@@ 0.D.5.3 Results
Actor learning rate 0.0001
Critic learning rate 0.001
@@
Clip ($\epsilon$) 0.1
Initial Standard Deviation 0.4
Standard Deviation Decay Rate 40000 steps
@@
Figure 21: Anomalies Penalty ($P_{M}$) by step for thirty Monte Carlo simulations (dashed lines) using simplified dynamics (red) and realistic dynamics (blue). Solid red and blue lines represent the average for simulations with simple and realistic dynamics, respectively.
Appendix 0.E Exploratory Data Analysis of StarLink open data
0.E.1 Residuals Analysis
@@ 0.E.2 Residual Scenarios
Figures 22 to 25 illustrate the residuals for these scenarios. The overlapping region provides the most reliable assessment of the propagator’s accuracy, as it is based on the most recent and accurate ephemeris data.
0.E.3 Figures and Observations
Figures 26 to 33 illustrate the satellites’ positional evolution and orbital trajectories over three days. These figures demonstrate the high accuracy of Orekit’s propagator when compared to Starlink’s ephemeris predictions, particularly in the overlapping regions.
Appendix 0.F Comparison of Orekit Propagation and Ephemerides
@@
0.F.1 Figures of Residuals and Orbits
Figures 39 to 42 present the propagated orbits compared to the ephemeris data. Oscillatory residuals were observed in the $x$, $y$, and $z$ components, increasing over time due to unmodeled perturbative forces and numerical integration errors.
Appendix 0.G Data and residual analysis

### v2

@@ E.4.3 Results
Experiment 1 10000 + $R_{E}$ km 0.1 $5.0^{\circ}$ $^{\circ} 10.010.0^{\circ}$ $10.0^{\circ}$
Experiment 2 15000 + $R_{E}$ km 0.5 $8.0^{\circ}$ $^{\circ} 10.010.0^{\circ}$ $90.0^{\circ}$
Experiment 3 5000 + $R_{E}$ km 0.001 $5.0^{\circ}$ $10.0^{\circ}$ $^{\circ} 180.0180.0^{\circ}$
@@
Figure 21: Reward per step, for each experiment. A reward of 0.0 indicates that the follower contains exactly the same orbital elements as the target, including anomaly.
@@
This experiment demonstrates that the agent learns an effective policy capable of generalizing thrust maneuvers to initial conditions it did not encounter during training, while also adapting to realistic orbital dynamics with non-conservative forces.
@@
E.4.4 Challenges and Future Improvements
Despite the agent demonstrating good generalization capabilities, it fails to perform maneuvers effectively when the target orbit’s inclination differs significantly from those encountered during training. Additionally, fuel consumption is not explicitly accounted for in the reward function, which negatively impacts the agent’s long-term strategy, often resulting in rapid fuel depletion.
Future improvements to this mission could include introducing additional penalties for fuel consumption in the reward function, training the agent with initial random orbits, and further exploring the limits of the agent’s strategy.
E.5 Collision Avoidance
E.5.1 Definition
@@
where $\hat{I}$, $\hat{J}$ and $\hat{K}$ are the unit vectors forming the orthogonal system. Similarly to the orbital plane encountered in a Keplerian orbit (as explained in B.1), given the assumption of negligible velocity uncertainty, only the $\hat{J}$ and $\hat{K}$ axes (defining the conjunction plane) are relevant for the analysis. This simplifies the problem to two dimensions by introducing a projection matrix $C=[\hat{J},\hat{K}$].
Under the assumption that uncertainties follow a Gaussian distribution, the distribution’s parameters are given by $\mu=C^{T}r_{\text{RTN}}$ and $\Sigma=C(\Sigma_{\text{T,RTN}}+\Sigma_{\text{C,RTN}})C^{T}$. The PoC is then calculated as the integral over a circle centered at the origin with radius $R=R_{T}+R_{C}$:
$\text{PoC}=\frac{1}{2\pi|\Sigma|^{\frac{1}{2}}}\int_{-R}^{R}\int_{-\sqrt{R^{2}-y^{2}}}^{\sqrt{R^{2}-y^{2}}}e^{-\frac{1}{2}\delta^{T}\Sigma^{-1}\delta}dzdy,$ (37)
@@
In this approach, we use a faster computation method compared to the traditional one, as outlined in [25]. According to NASA, a PoC value greater than $10^{-4}$ indicates a high collision risk, requiring maneuvers to address it. However, SpaceX places its threshold at $10^{-6}$, which is the one used in this experiment.
E.5.2 Environment Setup
Objective and Environment Characteristics.
@@ E.5.2 Environment Setup
Action Space.
As in the Hohmann maneuver mission (E.3) mission, apart from the polar representation of the thrust, the action space also contains the decision ($\delta$) to perform the given thrust: $a_{t}=(T,\theta,\phi,\delta)$. This action space is limited by $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}},\delta_{\text{max}})=(5,\pi,2\pi,1)$.
Observation Space.
The observation $o_{t}=(s_{t},M_{t},s^{\prime}_{t},M^{\prime}_{t},f_{t},P_{t})\in\mathbb{R}^{14}$ consists of the current equinoctial elements of the satellite ($s_{t}$) and the drifter ($s^{\prime}_{t}$), both bodies mean anomaly ($M_{t}$ and $M^{\prime}_{t}$), together with the current fuel mass ($f_{t}$) and PoC ($P_{t}$).
Reward Function.
@@
[…]
E.5.3 Results
The weights used to measure the deviation from the nominal orbit ($\Delta s_{t}$) were set to
$w=(10^{1},10^{-2},10^{-2},10^{-1},10^{-1}),$
and the weights assigned to each term in the reward function were $(\alpha_{1},\alpha_{2})=(10^{0},10^{-1})$. No extensive tuning was performed to find optimal parameter values.
For collision avoidance maneuvers, algorithms that work with discrete actions, such as DQN, work well, as shown by [13]. Because of this, we used both the DDQN (with the implementation shown in section E.1.3) and PPO (with the implementation shown in section E.1.1) algorithms. The training hyperparameters for both algorithms are shown in Table 12.
Table 12: Collision Avoidance: Training hyperparameters for PPO and DDQN.
PPO
@@ E.5.3 Results
DDQN
Parameter Value
Learning rate 0.00005
Epochs 1
Discount factor ($\gamma$) 0.95
$\tau$ 0.001
Memory Capacity 10 000
Initial $\epsilon$ 0.5
$\epsilon$ Decay Rate 1000 steps
$\epsilon$ Decay Amount 0.05
Minimum $\epsilon$ 0.05
Batch Size 256
Update Target Frequency 10 Online updates
Since DDQN only handles discrete actions, we chose to consider the maximum application of thrust in each of the directions in a 3D space, together with the non-action, as shown in Table 13.
Table 13: Possible actions in DDQN.
Action Thrust $(T,\theta,\phi,\delta)$ Direction
0 $(5,0,0,1)$ Forward
1 $(5,\pi/2,0,1)$ Left (radial out)
2 $(5,\pi,0,1)$ Behind
3 $(5,\pi/2,\pi,1)$ Right (radial in)
4 $(5,\pi/2,\pi/2,1)$ Up
5 $(5,\pi/2,3\pi/2,1)$ Down
6 $(0,0,0,0)$ Nothing
The training progression for both algorithms is illustrated in Figure 23, where PPO demonstrates more stable convergence compared to DQN. To evaluate the agent’s performance, we examine its decision-making behavior through key metrics such as the PoC, miss distance, and TCA. Furthermore, to assess the agent’s generalization capabilities, we activate all perturbative forces available in OrbitZoo, including higher-order gravitational harmonics, third-body perturbations from the Sun, Moon and Jupiter, SRP, and atmospheric drag. Importantly, the internal satellite simulator dynamics remain unchanged, preserving the realistic limitation that the agent operates under incomplete information about the environment.
Figure 23: Collision Avoidance: Cumulative reward per episode using DDQN and PPO.
The DDQN results (Figure 24) show that, under unrealistic dynamics, the agent effectively reduces the collision risk to a low level before the time of closest approach (TCA) by occasionally activating thrust. Under realistic dynamics, although the agent experiences a consistently low collision risk throughout the episode, it applies thrust almost continuously. A detailed examination of the orbital elements in Figure 25 reveals that, under the unrealistic dynamics encountered during training, the agent performs extremely well by maintaining the nominal orbit. In contrast, under realistic dynamics, the agent reduces its semi-major axis by approximately 40 km while keeping the eccentricity components close to those of the nominal orbit.
@@ 6 $(0,0,0,0)$ Nothing
Figure 24: Collision Avoidance using DDQN: Agent decision to apply thrust ($\delta>0.5$), Probability of Collision (POC), Miss Distance, and Time of Closest Approach (TCA) by step.
@@
Figure 25: Collision Avoidance using DDQN: Evolution of the semi-major axis (a) and eccentricity components ($e_{x}$ and $e_{y}$) of the orbit by step, using unrealistic and realistic dynamics.
In contrast to DDQN, the results from PPO (Figure 26) indicate that the agent learns a strategy of applying thrust early in the episode – precisely when the collision risk is highest. This early intervention effectively reduces the risk for the remainder of the episode. Notably, the agent’s behavior closely resembles what would be expected under realistic orbital dynamics, demonstrating the effectiveness of the learned strategy. A closer examination of the orbital elements in Figure 27 shows that the agent increases the semi-major axis by approximately 3 km while maintaining the eccentricity components near those of the nominal orbit.
Figure 26: Collision Avoidance using PPO: Agent decision to apply thrust ($\delta>0.5$), Probability of Collision (POC), Miss Distance, and Time of Closest Approach (TCA) by step.
Figure 27: Collision Avoidance using PPO: Evolution of the semi-major axis (a) and eccentricity components ($e_{x}$ and $e_{y}$) of the orbit by step, using unrealistic and realistic dynamics.
From these experiments, we conclude that although both agents learn effective policies under the unrealistic dynamics used during training, PPO demonstrates superior performance under realistic dynamics. This improvement is largely attributed to the continuous action space in PPO, which provides greater flexibility in executing maneuvers.
E.5.4 Challenges and Future Improvements
While both algorithms demonstrate some ability to generalize to realistic dynamics, they struggle with the inherent instability of the satellite simulation, particularly the sharp spikes in probability of collision (PoC) near the time of closest approach (TCA). Additionally, the agents have difficulty generalizing their strategies when initialized on orbits that differ from those encountered during training. Although DDQN is unable to fully generalize to realistic dynamics, it shows potential by maintaining consistent eccentricity components, suggesting a partial understanding of the objective.
Future improvements could include training with a randomized time horizon for maneuver completion, initializing agents from a broader range of orbits, and enhancing the observation space to include more informative features. These modifications would reduce uncertainty and support more effective decision-making. Additionally, providing more detailed information to each body’s internal simulation could further enable robust strategy learning. The DDQN architecture may also benefit from the integration of recurrent networks, as demonstrated by [13], which are capable of capturing temporal dependencies critical for decision-making in dynamic environments.
E.6 GEO Constellation
E.6.1 Definition
@@
There are various possible formations for constellations of satellites in geostationary orbit (GEO). In this work, we focus on a mission scenario where all agents learn to maintain equal angular separations along the orbit. Specifically, with $n$ satellites, each pair of adjacent satellites should be separated by an anomaly difference of $2\pi/n$ rad.
Given that the satellites share similar observations and are cooperating, we assume that agents can exchange knowledge after training, as discussed in Section C.5. In the context of PPO, this enables agents to share their critic networks after local training by aggregating and averaging model parameters – an operation that could be performed, for example, by a ground station. By forming a global critic, each agent gains a more comprehensive understanding of the environment’s overall state, even though it has only partial observations locally.
E.6.2 Environment Setup
Objective and Environment Characteristics.
We define an environment with $n=4$ agents in GEO, characterized by a semi-major axis of $a_{\text{GEO}}=35786+R_{E}=42164$ km, where $R_{E}$ is Earth’s equatorial radius, in a near perfect circular orbit. Each agent is a satellite with $m_{0}=250$ kg, 50 kg of which is fuel, and equipped with a thrust with $I_{sp}=3100$ s.
Each episode contains 500 steps with step sizes of 360 s (approximately 2 revolutions/days), and the dynamics are purely based on Newtonian attraction. At the start of each episode ($t=0$), agents are initialized with random mean anomalies ($M_{0}^{i},i=\{1,2,...,n\}$), representing their positions along the orbit. The objective for the agents is to learn to distribute themselves uniformly around the orbit, maintaining angular separations of $2\pi/n$ radians (90 degrees for $n=4$).
Action Space.
@@ E.6.2 Environment Setup
Observation Space.
The observation space is also reduced. Each agent observation $o_{t}=(a_{t},e_{xt},e_{yt},f_{t},M_{t}^{1},M_{t}^{2},M_{t}^{3},M_{t}^{4})$ receives its current semi-major axis, eccentricity components, remaining fuel, and the anomalies of all agents in the environment.
Reward Function.
To calculate the collective penalty related to the difference in anomalies ($P_{M}$), we first normalize the anomalies to be within $0,2\pi]$, followed by:
@@
E.6.3 Results
The agents were trained both independently (without communication) and with a centralized critic using the Federated Averaging (FedAvg) algorithm, with the training hyperparameters provided in Table 14.
For training, a combination of parameters that ended up performing rather well was $(\alpha_{1},\alpha_{2},\alpha_{3})=(10^{-8},10^{1},10^{-2})$. The training evolution using these parameters for both experiments is shown at Figure 28.
Figure 28: GEO Constellation: Average Cumulative Reward per episode, using independent and federated learning.
For evaluation, we tested the generalization capability of the policies by introducing realistic dynamics, which include harmonic gravity fields, third body forces, SRP, and drag. To evaluate the performance of both constellations, we conducted Monte Carlo simulations analyzing the evolution of the anomaly penalty, as illustrated in Figure 29.
Figure 29: Average Anomaly Penalty ($P_{M}$) per step over thirty Monte Carlo simulations with realistic dynamics, comparing independent and federated learning. In these simulations, initial anomalies are random but equal for both federated and independent constellations.
The results show that agents are capable of reducing the collective anomaly penalty in both scenarios. However, the rate and effectiveness of this reduction depend on the initial constellation configuration. Federated learning demonstrates a faster decrease in the anomaly penalty, highlighting its potential advantage over independent learning. Through the OrbitZoo interface, shown in Figure 2, it is evident that the agents are successfully learning the intended objective, albeit with some imperfections and rather slowly.
In conclusion, although the agents were initially trained independently without sharing information, they successfully learned to maintain equal spacing around the GEO orbit – even under realistic dynamics not encountered during training. The federated approach, however, demonstrated clear advantages, enabling more stable training and a better understanding of the global objective in a cooperative setting.
E.6.4 Challenges and Future Improvements
In some initial configurations – particularly those where agents begin well separated – the anomaly penalty fails to decrease significantly. This limitation appears in both the independent and federated constellations. Addressing it may require tuning hyperparameters more effectively, including adjustments to the reward function or exploring alternative RL architectures. Another challenge stems from the current reward setup: the agents must balance reducing the anomaly penalty with maintaining the nominal altitude, which slows the rate of improvement.
Future work includes implementing vertical federated learning (VFL) using split learning, as the FedAvg approach can become unstable due to its simplistic averaging mechanism. Another key direction is scaling to larger constellations to better reflect realistic operational scenarios.
@@ E.6.4 Challenges and Future Improvements
Actor learning rate 0.00001
Critic learning rate 0.0001
@@
Clip ($\epsilon$) 0.2
Initial Standard Deviation 0.5
Standard Deviation Decay Rate 10000 steps
@@
Appendix F Exploratory Data Analysis of StarLink open data
F.1 Residuals Analysis
@@ F.2 Residual Scenarios
Figures 30 to 33 illustrate the residuals for these scenarios. The overlapping region provides the most reliable assessment of the propagator’s accuracy, as it is based on the most recent and accurate ephemeris data.
F.3 Figures and Observations
Figures 34 to 41 illustrate the satellites’ positional evolution and orbital trajectories over three days. These figures demonstrate the high accuracy of Orekit’s propagator when compared to Starlink’s ephemeris predictions, particularly in the overlapping regions.
Appendix G Comparison of Orekit Propagation and Ephemerides
@@
G.1 Figures of Residuals and Orbits
Figures 47 to 50 present the propagated orbits compared to the ephemeris data. Oscillatory residuals were observed in the $x$, $y$, and $z$ components, increasing over time due to unmodeled perturbative forces and numerical integration errors.
Appendix H Data and residual analysis